# Analytical approach for calculating shadow of dynamical black hole

### v1

@@ Abstract
A real astrophysical black hole, surrounded by an accretion disc, experiences time-dependent mass variation due to accretion, impacting observable features like the shadow. In this work, we analyzed the dynamical shadow by extending the concept of the photon sphere to time-dependent spacetimes, incorporating induced, centrifugal, and relativistic forces. We demonstrated that the shadow grows when light loses energy and shrinks when it gains energy, with a surprising result that the shadow remains static when the photon sphere radius coincides with the NEC horizon radius. Our method, assuming a small accretion rate, provides a foundation for studying dynamical black holes, offering insights into processes like gravitational collapse, which could make shadow evolution detectable and reveal details about black hole formation.
@@ I Introduction
The shadow of a black hole has been proposed as a powerful tool for cosmological studies, serving as a “cosmological ruler” to measure and test various phenomena Tsupko et al. (2020); Vagnozzi et al. (2023); Chen et al. (2022); Allahyari et al. (2020). Through its observation, different black hole models can be rigorously tested Kuang and Övgün (2022); Kuang et al. (2022); Meng et al. (2022); Wang et al. (2023); Vagnozzi et al. (2023); Abdujabbarov et al. (2016a); Atamurotov et al. (2013); Abdujabbarov et al. (2015); Atamurotov et al. (2021); Gralla et al. (2019); Cunha and Herdeiro (2018); Afrin et al. (2021); Abdujabbarov et al. (2016b); Younsi et al. (2016); Konoplya and Zinhailo (2020); Konoplya (2019); Kumar Walia (2024); Rodríguez et al. (2024); Wang et al. (2017); Pantig and Övgün (2023, 2023); Pantig et al. (2022, 2023); Kumar and Ghosh (2020a, b); Okyay and Övgün (2022); Heidari et al. (2024); Lambiase et al. (2024a, b); Övgün et al. (2024); Karshiboev et al. (2024); Puliçe et al. (2023); Kumaran and Övgün (2023); Chakhchi et al. (2025, 2024); Belhaj et al. (2021), and intriguingly, even singularities—regions of spacetime where densities become infinite—can cast shadows Shaikh et al. (2019); Heydarzade and Vertogradov (2024).
@@
This paper is structured as follows: Section 2 provides an in-depth discussion of geodesic motion in both static and dynamical spacetimes, establishing a connection between the energy of a particle and the generalized flux. Section 3 is dedicated to calculating the dynamical radius of the photon sphere and the black hole shadow, incorporating the effects of a changing spacetime geometry. Section 4 concludes the research, summarizing the key findings and their implications.
[…]
$\dot{M}$ $=$ $\frac{\partial M}{\partial v},$
$M^{\prime}$ $=$ $\frac{\partial M}{\partial r}.$
@@ II Geodesic equations
We begin by exploring the foundational concepts of geodesic motion. Initially, we describe the motion of geodesics within a general spherically symmetric spacetime, emphasizing the fundamental principles governing such systems. Subsequently, we extend our analysis to the case of dynamical spacetimes, examining how the introduction of time dependence alters the behavior of geodesics.
By comparing the two models, we identify both their differences and similarities, shedding light on how dynamical effects influence geodesic motion.
@@ II.1 Static case
The general line element in Eddington-Finkelstein coordinates $\{v,r,\theta,\varphi\}$ for a spherically symmetric spacetime can be expressed as:
$ds^{2}$ $=$ $-fdv^{2}+2\varepsilon dvdr+r^{2}d\Omega^{2},$
$f(r)$ $=$ $1-\frac{2M(r)}{r}.$ (1)
where:
  • $M(r)$ is the mass function that characterizes the gravitational field as a function of the radial coordinate r.
  • $\varepsilon=\pm 1$ is a parameter, often used to distinguish between different coordinate conventions, which has no impact on geodesic motion in the static case.
  • $d\Omega^{2}=d\theta^{2}+\sin^{2}\theta d\varphi^{2}$ is the metric on the unit two-sphere, representing the angular part of the spacetime in spherical coordinates.
This form is particularly useful for studying geodesic motion because it separates the radial, temporal, and angular components of the spacetime, making it easier to analyze the dynamics in both static and dynamical cases a spacetime (II.1) is spherically-symmetric and static. The spacetime admits two Killing vectors, reflecting its symmetries under time translation and rotation. These symmetries result in two conserved quantities along geodesics: Energy per unit mass $(E)$, associated with the time-translation invariance. Angular momentum per unit mass $(L)$, stemming from rotational invariance.
$E$ $=$ $f\frac{dt}{d\lambda},$
$L$ $=$ $r^{2}\sin^{2}\theta\frac{d\varphi}{d\lambda},$ (2)
where the $\lambda$ is an affine parameter. In a spherically symmetric spacetime, the motion of particles and light can always be confined to a single plane due to the rotational symmetry. Without loss of generality, we choose the equatorial plane, where $\theta=\frac{\pi}{2}$. In this plane, the angular derivative terms vanish, i.e., $\frac{d^{2}\theta}{d\lambda^{2}}=\frac{d\theta}{d\lambda}=0$. This simplifies the analysis to motion in the equatorial plane. Using the geodesic condition $g_{ik}u^{i}u^{k}=\delta$, one finds
$\left(\frac{dr}{d\lambda}\right)^{2}$ $+$ $V_{eff}=0,$
$V_{eff}$ $=$ $f\left(\frac{L^{2}}{r^{2}}-\delta\right)-E^{2}.$ (3)
The parameter $\delta=0,\pm 1$ in the geodesic equation identifies the type of trajectory being considered: null geodesics $(\delta=0)$ describe the motion of light, timelike geodesics $(\delta=-1)$ correspond to the motion of massive particles, and spacelike geodesics $(\delta=+1)$ represent less common, hypothetical paths. The effective potential, $V_{\text{eff}}$, governs the motion and shapes the trajectory of the particle or photon, encapsulating the combined influence of gravitational and centrifugal forces. For circular orbits, where the radial coordinate r remains constant, two conditions must be met: the effective potential must vanish $(V_{\text{eff}}=0)$, and its first derivative with respect to r must also be zero $(\frac{dV_{\text{eff}}}{dr}=0)$. Additionally, the stability of the circular orbit is determined by the second derivative of the effective potential $(\frac{d^{2}V_{\text{eff}}}{dr^{2}})$. If $\frac{d^{2}V_{\text{eff}}}{dr^{2}}>0$, the orbit is stable, while $\frac{d^{2}V_{\text{eff}}}{dr^{2}}<0$ indicates an unstable orbit.
$\frac{d^{2}r}{d\lambda^{2}}=-\frac{1}{2}V^{\prime}_{eff},$ (4)
which leads to
$\frac{d^{2}r}{d\lambda^{2}}=\delta\frac{M-M^{\prime}r}{r^{2}}+\left(M^{\prime}r-3M\right)\frac{L^{2}}{r^{4}}+\frac{L^{2}}{r^{3}},$ (5)
where $M{\prime}$ is the derivative of the mass function with respect to $r$, and $M$ is the mass function itself. $L$ is the angular momentum per unit mass.
We identify three key terms that contribute to the effective forces governing motion in the spacetime:
  • Newtonian Attractive Force $(a_{n})$: This term represents the classical gravitational attraction, given by: $a_{n}$
    $a_{n}=\frac{M^{\prime}r-M}{r^{2}}.$ (6)
  • Centrifugal Repulsive Force $a_{c}$: This term accounts for the repulsive effect of angular momentum, expressed as:
    $a_{c}=\frac{L^{2}}{r^{3}}.$ (7)
  • General Relativistic Corrections $a_{g}$: This term arises from general relativistic effects and modifies the overall dynamics of the system. Its specific form will depend on the spacetime under consideration and encapsulates deviations from classical mechanics due to the curvature of spacetime.
    $a_{g}=\frac{M^{\prime}r-3M}{r^{4}}L^{2}.$ (8)
Hence, the geodesic equation (5), incorporating the identified force terms, can be expressed as:
$\frac{d^{2}r}{d\lambda^{2}}=-\delta a_{n}+a_{c}+a_{g}.$ (9)
It is important to note that the Newtonian attractive force, $a_{n}$, is present only for timelike or spacelike geodesic motions. The radius $r_{ph}$ is the radius of the photon sphere if the following conditions are satisfied:Claudel et al. (2001)
  • The effective radial acceleration is zero, indicating equilibrium:
    $V_{eff}\left(r_{ph}\right)=0,~{}~{}a_{g}\left(r_{ph}\right)+a_{c}\left(r_{ph}\right)=0.$ (10)
  • The second condition for the photon sphere radius, $r_{ph}$, follows from the requirement that the derivative of the effective potential vanishes at that radius, i.e.,$V^{\prime}_{eff}\left(r_{ph}\right)=0$. This ensures that the forces at $r_{ph}$ are in equilibrium.
Therefore, we can reformulate the definition of the photon sphere as follows:
A radius of a photon sphere is the radius at which the centrifugal repulsive force and general relativistic corrections are balanced.
This revised definition provides a clear physical interpretation and establishes a basis for extending the concept to a dynamical photon sphere, which will be discussed further below.
@@ II.2 Dynamical case
We consider a spherically symmetric dynamical spacetime in Eddington-Finkelstein coordinates $v,r,\theta,\varphi$ described by the following line element:
$ds^{2}=-f(v,r)dv^{2}+2\varepsilon dvdr+r^{2}d\Omega^{2},$ (11)
and the function $f(v,r)$ represents the lapse function and is given by:
$f(v,r)=1-\frac{2M(v,r)}{r},$ (12)
with $M(v,r)$ as the mass function, which depends on both the Eddington time coordinate $v$ and the radial coordinate $r$. The parameter $\varepsilon=\pm 1$ distinguishes between ingoing $(\varepsilon=+1)$ and outgoing $(\varepsilon=-1)$ radiation. In a spherically symmetric spacetime, the symmetry allows us to simplify the analysis of geodesic motion by restricting it to the equatorial plane, where $\theta=\frac{\pi}{2}$.
To describe the geodesic equation, we start with the action $I$, which governs the motion of a particle or photon in a curved spacetime. The action is given by:
[…]
The resulting equations of motion for each coordinate are as follows:
$\frac{d\varphi}{d\lambda}=\frac{L}{r^{2}},$ (14)
$-\frac{1}{2}f^{\prime}\left(\frac{dv}{d\lambda}\right)^{2}+r\left(\frac{d\varphi}{d\lambda}\right)^{2}-\varepsilon\frac{d^{2}v}{d\lambda^{2}}=0,$ (15)
$\varepsilon\frac{d^{2}r}{d\lambda^{2}}=\frac{1}{2}\dot{f}\left(\frac{dv}{d\lambda}\right)^{2}+f\frac{d^{2}v}{d\lambda^{2}}+f^{\prime}\frac{dv}{d\lambda}\frac{dr}{d\lambda}.$ (16)
In the equations of motion derived from the action, $\lambda$ represents the affine parameter along the geodesic, which parameterizes the trajectory of a particle or photon. The conserved angular momentum per unit mass is denoted by $L$, a quantity arising from the rotational symmetry of the spacetime. A dot $\dot{}$ refers to the partial derivative with respect to the Eddington-Finkelstein time coordinate $v$. A dash ($\prime$) indicates the partial derivative with respect to the radial coordinate $r$.
Expressing $f\frac{d^{2}v}{d\lambda^{2}}$ from (15) by using (14), we obtain
$f\frac{d^{2}v}{d\lambda^{2}}=\varepsilon f\frac{L^{2}}{r^{3}}-\frac{1}{2}\varepsilon ff^{\prime}\left(\frac{dv}{d\lambda}\right)^{2}.$ (17)
The null geodesic condition $g_{ik}\frac{dx^{i}}{d\lambda}\frac{dx^{k}}{d\lambda}=0$ gives
$f^{\prime}\frac{dv}{d\lambda}\frac{dr}{d\lambda}=\frac{1}{2}\varepsilon ff^{\prime}-\frac{1}{2}\varepsilon f^{\prime}\frac{L^{2}}{r^{2}}\left(\frac{dv}{d\lambda}\right)^{2}.$ (18)
Now, substituting (17) and (18) into (16), we finaly obtain
$\frac{d^{2}r}{d\lambda^{2}}=\frac{1}{2}\varepsilon\dot{f}\left(\frac{dv}{d\lambda}\right)^{2}-\frac{1}{2}f^{\prime}\frac{L^{2}}{r^{2}}+f\frac{L^{2}}{r^{3}}.$ (19)
By substituting (12) we arrive at general form of geodesic equation
$\frac{d^{2}r}{d\lambda^{2}}=-\varepsilon\frac{\dot{M}}{r}\left(\frac{dv}{d\lambda}\right)^{2}+\left(M^{\prime}r-3M\right)\frac{L^{2}}{r^{4}}+\frac{L^{2}}{r^{3}}.$ (20)
The first term on the right hand-side of (20) represents a non-Newtonian gravitational force arising from the dynamics of the black hole, specifically its radiation or accretion processes, which alter the mass function $M(v,r)$ over time. This term captures the energy exchange between the black hole and its surroundings, distinguishing it from the static Newtonian force. The second term accounts for relativistic corrections predicted by general relativity, which modify the gravitational force significantly in strong-field regions, such as near the event horizon or the photon sphere. Lastly, the third term corresponds to the repulsive centrifugal force resulting from the angular momentum of the particle or photon, which opposes the inward pull of gravity and establishes equilibrium in circular motion Vertogradov and Kudryavcev (2023).
The generalized total apparent flux, $\Lambda$, associated with the black hole’s accretion rate, can be defined as Lindquist et al. (1965):
$\Lambda=\varepsilon\dot{M}\left(\frac{dv}{d\lambda}\right)^{2}\,,$ (21)
We refer to the first term on the right-hand side as the induced acceleration, denoted by $a_{i}\equiv-\frac{\Lambda}{r}$. Additionally, we denote the general relativistic correction and the repulsive centrifugal force as $a_{g}$ and $a_{c}$, respectively. With these definitions, Eq. (20) can be rewritten as:
$\frac{d^{2}r}{d\lambda^{2}}=a_{i}(v,r)+a_{g}(v,r)+a_{c}(v,r).$ (22)
In dynamical spacetimes, it is a well-known fact that the energy per unit mass of a particle is not conserved along null geodesics due to the time-dependent nature of the spacetime. However, one can still define this energy using the Lagrangian formalism.
$2\mathcal{L}=g_{ik}\frac{dx^{i}}{d\lambda}\frac{dx^{k}}{d\lambda},$ (23)
as follows
$E(v)=-\frac{\partial\mathcal{L}}{\partial\dot{v}}=f\frac{dv}{d\lambda}-\frac{dr}{d\lambda}.$ (24)
By using (14), (24) and null geodesic equation condition $g_{ik}u^{i}u^{k}=0$, one obtains
$\left(\frac{dr}{d\lambda}\right)^{2}+V_{eff}(v,r)=0,$
$V_{eff}(v,r)=f\frac{L^{2}}{r^{2}}-E^{2}(v).$ (25)
To express the energy as a function of the total generalized flux $\Lambda$, we start by differentiating the energy expression with respect to the affine parameter $\Lambda$. From the definition of energy (II.2):
$02\frac{dr}{d\lambda}\left(\frac{d^{2}r}{d\lambda^{2}}+\frac{1}{2}\frac{\partial V_{eff}}{\partial r}\right)+\frac{\partial V_{eff}}{\partial v}\frac{dv}{d\lambda}=$ 0 . (26)
However
$\frac{1}{2}\frac{\partial v_{eff}}{\partial r}=a_{g}-a_{c}.$ (27)
By using (20) we arrive at
$-2\frac{\Lambda}{r}\frac{dr}{d\lambda}+\frac{\partial V_{eff}}{\partial v}\frac{dv}{d\lambda}=0.$ (28)
Remembering the definition of (21) and using the fact that
$\frac{\partial V_{eff}}{\partial v}$ $=$ $-\frac{2\dot{M}L^{2}}{r^{3}}-2E\dot{E},$
$2\frac{dr}{d\lambda}\frac{dv}{d\lambda}$ $=$ $f\left(\frac{dv}{d\lambda}\right)^{2}-\frac{L^{2}}{r^{2}},$ (29)

### v2

@@ Abstract
We develop a compact and transparent framework for photon dynamics and shadow formation in slowly evolving, spherically symmetric spacetimes. Starting from the Eddington-Finkelstein action, we derive a force-decomposed radial equation in which the radial acceleration splits into an induced term sourced by mass variation, a centrifugal term, and a purely general-relativistic correction. A key result is a gauge-invariant energy-flux relation, $d(E^{2})/dv=-\Lambda/r$, with $\Lambda\equiv\varepsilon\,\dot{M}\,\dot{v}^{2}$, which controls how time dependence modifies the canonical energy of null geodesics. In the adiabatic regime we obtain an explicit first-order shift of the photon-sphere radius, $r_{ph}(v)=r_{0}-a_{i}/(a_{g}^{\prime}+a_{c}^{\prime})$, and connect it to the observable shadow through the evolving critical impact parameter, $b_{\rm crit}(v)=\sqrt{r_{ph}(v)^{2}/f(r_{ph}(v))}$. For Vaidya spacetimes this predicts that accretion ($\dot{M}>0$) expands the photon sphere and increases the shadow angle, whereas mass loss has the opposite effect. Our formulation refines classic force-balance ideas to dynamical settings, provides a constructive link to time-dependent photon surfaces, and yields simple, observer-ready expressions for the evolution of the shadow. The framework offers a baseline for confronting time-variable horizon-scale imaging with dynamical inflow/outflow models.
@@ I Introduction
The shadow of a black hole has been proposed as a powerful tool for cosmological studies, serving as a “cosmological ruler” to measure and test various phenomena Tsupko et al. (2020); Vagnozzi et al. (2023); Chen et al. (2022); Allahyari et al. (2020). Through its observation, different black hole models can be rigorously tested Tsukamoto et al. (2014); Tsukamoto (2018); Kuang and Övgün (2022); Kuang et al. (2022); Meng et al. (2022); Wang et al. (2023); Vagnozzi et al. (2023); Abdujabbarov et al. (2016a); Atamurotov et al. (2013); Abdujabbarov et al. (2015); Atamurotov et al. (2021); Gralla et al. (2019); Cunha and Herdeiro (2018); Afrin et al. (2021); Abdujabbarov et al. (2016b); Younsi et al. (2016); Konoplya and Zinhailo (2020); Konoplya (2019); Kumar Walia (2024); Rodríguez et al. (2024); Wang et al. (2017); Pantig and Övgün (2023, 2023); Pantig et al. (2022, 2023); Kumar and Ghosh (2020a, b); Okyay and Övgün (2022); Heidari et al. (2024); Lambiase et al. (2024a, b); Övgün et al. (2024); Karshiboev et al. (2024); Puliçe et al. (2023); Kumaran and Övgün (2023); Chakhchi et al. (2025, 2024); Belhaj et al. (2021); Liu et al. (2024a, b), and intriguingly, even singularities—regions of spacetime where densities become infinite—can cast shadows Shaikh et al. (2019); Heydarzade and Vertogradov (2024).
@@
This paper is organized as follows. In Sec. II we develop the geodesic dynamics in spherically symmetric spacetimes, first in the static case and then in the dynamical case, and we derive a clean relation between the particle energy and a generalized flux that encodes accretion/radiation. Section III applies these results to obtain the adiabatic evolution of the photon-sphere radius and of the black-hole shadow. We summarize our main conclusions and outlook in Sec. IV.
Throughout we use geometrized units with $G=c=1$ and the metric signature $\{-+++\}$. For the (generally) spacetime-dependent mass function $M(v,r)$ we denote
$\dot{M}\equiv\frac{\partial M}{\partial v},\qquad M^{\prime}\equiv\frac{\partial M}{\partial r}.$ (1)
@@ II Geodesic equations
We begin by laying out the kinematics of null geodesics in spherically symmetric geometries. Our aim is twofold: (i) to isolate the force-like terms that control radial acceleration and (ii) to identify how time dependence modifies the conserved energy and thus the observable shadow.
@@ II.1 Static case
Ingoing Eddington-Finkelstein (EF) coordinates $\{v,r,\theta,\varphi\}$ are regular across future horizons and align the time coordinate with ingoing null rays. For a static, spherically symmetric line element we use
$ds^{2}=-f(r)\,dv^{2}+2\,dv\,dr+r^{2}d\Omega^{2},\qquad f(r)=1-\frac{2M(r)}{r}.$ (2)
EF time $v$ is a Killing parameter here, so the definition of energy below is equivalent to the standard Schwarzschild definition and avoids coordinate singularities at the horizon. The Killing vector $\partial_{v}$ yields a conserved energy
$E=-p_{v}=f(r)\,\frac{dv}{d\lambda}-\frac{dr}{d\lambda},$ (3)
while SO(3) symmetry implies the conserved angular momentum
$L=r^{2}\sin^{2}\theta\,\frac{d\varphi}{d\lambda}.$ (4)
Motion can be restricted to the equatorial plane $\theta=\pi/2$ without loss of generality. With $g_{ik}u^{i}u^{k}=\delta$ ($\delta=0,-1,+1$ for null/timelike/spacelike), radial motion obeys
$\Big(\frac{dr}{d\lambda}\Big)^{2}+V_{\rm eff}(r;E,L,\delta)=0,\qquad V_{\rm eff}=f(r)\Big(\frac{L^{2}}{r^{2}}-\delta\Big)-E^{2}.$ (5)
For null geodesics ($\delta=0$), turning points occur where $V_{\rm eff}=0$. Circular orbits require $V_{\rm eff}=V^{\prime}_{\rm eff}=0$. Differentiating $V_{\rm eff}$ gives
$\frac{d^{2}r}{d\lambda^{2}}=\delta\,\frac{M-M^{\prime}r}{r^{2}}+\Big(M^{\prime}r-3M\Big)\frac{L^{2}}{r^{4}}+\frac{L^{2}}{r^{3}}.$ (6)
It is natural to define
$a_{n}\equiv\frac{M^{\prime}r-M}{r^{2}},\qquad a_{c}\equiv\frac{L^{2}}{r^{3}},\qquad a_{g}\equiv\frac{(M^{\prime}r-3M)L^{2}}{r^{4}},$ (7)
so that $\ddot{r}=-\delta\,a_{n}+a_{c}+a_{g}$. $a_{c}$ is the familiar centrifugal term; $a_{g}$ encodes purely GR curvature corrections and flips sign at $r=3M$ in Schwarzschild; $a_{n}$ is a Newtonian-like term present only for $\delta\neq 0$. Null circular orbits satisfy
$a_{g}(r_{ph})+a_{c}(r_{ph})=0,$ (8)
i.e. the GR correction exactly cancels the centrifugal term. In Schwarzschild this gives $r_{ph}=3M$, in line with the classic picture of centrifugal-force reversal near strong curvature.
@@ II.2 Dynamical case
We now allow for explicit time dependence through $M(v,r)$ and keep an $\varepsilon=\pm 1$ flag to cover both ingoing and outgoing EF charts:
$ds^{2}=-f(v,r)\,dv^{2}+2\varepsilon\,dv\,dr+r^{2}d\Omega^{2},\qquad f(v,r)=1-\frac{2M(v,r)}{r}.$ (9)
Spherical symmetry is retained, so $L$ remains conserved; however, $E$ is generically not conserved. From
$I=\tfrac{1}{2}\int\!\Big[-f(v,r)\,\dot{v}^{2}+2\varepsilon\,\dot{v}\,\dot{r}+r^{2}\dot{\varphi}^{2}\Big]\,d\lambda,$ (10)